Question 4

A manufacturer of car batteries claims that the life of his batteries is approximately normally distributed with a mean life of 7 years and a standard deviation equal to 2 year. What is the probability that a battery last between 1 and 13 years?

a.

p = 0.9974

b.

p =0.74

c.

p = 0.99

d.

p = -0.9974
Answer #1

Ans. Let X be life of battery, and X has N( 7, 4)

p = P( 1 < X < 13)

= P[(1-7)/2 < (X-7)/2 < (13-7)/2]

= P (-3 < Z < 3)

= 2*P(0 < Z < 3), since Z is symmetrical distribution abuta z =0

=2*0.4987

=9974

Thus the correct answer is (a)